# literature.bib

### v2

@@ 1 Introduction
The combination of bilevel optimization and uncertainty is a topic that receives increasing attention currently. In a practical setting in which leader and follower are actually two distinct decision makers, it seems to be very natural to adopt some concept of optimization under uncertainty and assume that, e.g., the leader is uncertain about some parameters of the follower’s problem. A very recent and comprehensive overview of the area of bilevel optimization under uncertainty is given by Beck, Ljubić, and Schmidt in the survey [beckljubicschmidt2023]. It covers both stochastic and robust approaches to model uncertainty, and it classifies various ways of where and how uncertainty can come into play in a bilevel optimization problem. In this article, we focus on uncertainty in the parameters of the follower’s objective function from the leader’s perspective. The situation can be illustrated by considering three decision makers: The leader first chooses a solution, then an adversary (of the leader) chooses an objective function for the follower, and the follower finally computes an optimal solution according to this objective function and depending on the leader’s choice. The leader’s aim is to optimize her own objective value, which depends on the follower’s choice, while the adversary has the opposite objective. Note that defender-attacker-defender games present a similar three-level structure as these robust bilevel optimization problems; see, e.g., [brownetal2006]. However, there is only one objective function in these problems, as the same decision maker acts on the first and on the third level. Moreover, the attack does usually not affect the objective function, but a combinatorial structure concerning the constraints.
@@
A related robust bilevel optimization problem, namely the Robust Bilevel Continuous Knapsack Problem, together with its complexity for different types of uncertainty sets, has been studied in [buchheimhenke2022]. A part of the current work generalizes and extends the results of [buchheimhenke2022], with a stronger emphasis on combinatorial underlying problems. The (Continuous) Bilevel Selection Problem with disjoint item sets is more general than the Bilevel Continuous Knapsack Problem in one aspect, but structurally simpler than it in another; see Section 2.2 for more details. For most of the robust settings we consider, we obtain similar results as in [buchheimhenke2022], but we will also see some differences. In particular, solving the Robust Continuous Bilevel Selection Problem under discrete uncorrelated uncertainty requires some new ideas, which also lead to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem; see Section 6.2.3. The latter result answers an open question stated in [buchheimhenke2022].
@@
This article is organized as follows. In Section 2, we introduce the Bilevel Selection Problem and discuss several variants of the problem, some of which are distinguished later on. We show how the Bilevel Selection Problem can be solved in polynomial time in Section 2.3. The robust problem version is introduced in Section 3. We then present algorithms for the adversary’s problem, for different types of uncertainty sets, in Section 4, before turning to the robust leader’s problem in Section 5. Here we give polynomial-time algorithms for the simpler case of disjoint item sets in Section 5.1 and show our hardness and algorithmic results for the general case in Section 5.2. Moreover, we give some insights regarding problem variants with continuous variables in Section 6. This includes an example of a bilevel optimization problem whose optimal value may not be attained (Example 25) and some structural properties in Section 6.1 and in the beginning of Section 6.2. We then study the Robust Continuous Bilevel Selection Problem for different types of uncertainty sets. A new approach developed here is finally extended to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty in Section 6.2.3.1. Section 7 concludes.
@@ 2.1 Problem Definition
Note that we do not require the sets ${\mathcal{E}_{l}}$ and ${\mathcal{E}_{f}}$ to be disjoint in general. However, assuming the sets to be disjoint leads to an important special case, actually simplifying the problem in some sense. We will see in Section 2.3 that our polynomial-time algorithm works for the special case and for the general case; however, it is more intuitive in the case of disjoint item sets. When it comes to the robust problem version, the special case with disjoint item sets seems to be significantly easier than the general case, as we discuss in Sections 5.1 and 5.2.
We denote the cardinalities of the item sets by ${n_{l}}=\lvert{\mathcal{E}_{l}}\rvert$, ${n_{f}}=\lvert{\mathcal{E}_{f}}\rvert$, and $n=\lvert{\mathcal{E}_{l}}\cup{\mathcal{E}_{f}}\rvert$ throughout this article.
@@
In fact, from the leader’s perspective, only this order of the follower’s items is relevant, but not the precise follower’s item costs. Accordingly, when constructing an instance, we will sometimes only define a follower’s greedy order instead of a function $d$; see, e.g., Theorem 15 and Examples 19 and 20. Observe that, given a follower’s greedy order, an appropriate function $d$ can easily be constructed in polynomial time, e.g., by setting its values to subsequent natural numbers in this order.
@@ 2.2 Problem Variants and Relation to Knapsack Problems
In this section, we introduce and discuss several assumptions and modifications that can be made for the Bilevel Selection Problem as we defined it in Section 2.1. Some of the resulting problem variants will be distinguished later on; in particular, this concerns the setting of disjoint leader’s and follower’s item sets and the general one in Section 5, and the problem variants with continuous variables in Section 6. For other variants mentioned in the following, we indicate how they are related to each other and to other problems that have been studied, in order to provide a broader picture of the Bilevel Selection Problem.
@@ 2.2.1 Item Sets
As mentioned in Section 2.1, we will mainly distinguish the special case of the Bilevel Selection Problem where ${\mathcal{E}_{l}}\cap{\mathcal{E}_{f}}=\emptyset$ from the general one. This special case of the Bilevel Selection Problem can be considered to be a special case of the Bilevel Knapsack Problem that has been studied in [mansietal2012, capraraetal2014]. In fact, the polynomial-time algorithm presented for the Bilevel Selection Problem in Section 2.3, in case of disjoint item sets, can be seen as a special case of the pseudopolynomial-time approaches in [mansietal2012, capraraetal2014]. However, our algorithm also works for the general Bilevel Selection Problem, where the situation is more involved.
Moreover, one could think of a simpler bilevel variant of the Selection Problem in which only the follower selects items in the sense of a single-level Selection Problem, while the leader just decides on the number $b$ of items to select and gets some value (which might be positive or negative) from each selected item. This problem variant can be seen as a special case of the Bilevel Knapsack Problem in which the leader provides the capacity for the follower’s knapsack problem. This problem has been studied in [demperichter2000, capraraetal2014], and the continuous version of it is the underlying Bilevel Continuous Knapsack Problem in [buchheimhenke2022, buchheimhenkeirmai2022]. In fact, this bilevel selection problem is equivalent to a special case of the Bilevel Selection Problem as defined above, with disjoint item sets: Consider ${\mathcal{E}_{l}}$ and ${\mathcal{E}_{f}}$ to be disjoint and the leader’s costs of all leader’s items to be equal to zero. Then, by selecting some subset $X\subseteq{\mathcal{E}_{l}}$ of her items, the leader just determines the capacity $b-\lvert X\rvert$ for the follower and pays her item costs of all items selected by the follower. Therefore, the algorithmic results we obtain in Sections 2.3, 5.1, and 6 can also be applied to this bilevel selection problem, and in fact are similar to the ones in [buchheimhenke2022]. In Section 6.2.3.1, we extend our new algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncorrelated uncertainty to the Robust Bilevel Continuous Knapsack Problem.
@@ 2.2.3 Maximization
The Selection Problem and also the Bilevel Selection Problem are defined as minimization problems according to Section 2.1, but could analogously be defined as maximization problems, i.e., the decision makers could maximize the total value of their selection of items instead of minimizing the total costs. Then the problems can be more directly seen as special cases of the classical Knapsack Problem and a bilevel variant of it; see also Section 2.2.1. In Section 6.2.3.1, we transfer a new algorithm from the Robust Continuous Bilevel Selection Problem to the Robust Bilevel Continuous Knapsack Problem, which also requires to adjust some details regarding the direction of optimization.
@@ 5.1 The Special Case of Disjoint Item Sets
The Robust Bilevel Selection Problem with ${\mathcal{E}_{l}}\cap{\mathcal{E}_{f}}=\emptyset$, and with any type of uncertainty set, can be solved in time $O(nR(I))$, where, for an instance $I$ of the Robust Bilevel Selection Problem, $R(I)$ is an upper bound on the running time for solving the corresponding adversary’s problem with any fixed leader’s solution given.
@@
The leader enumerates all feasible values of ${b_{l}}$ that determine the capacity used by herself. These are clearly the integer values in the same range as in Algorithm 1, and their number is $O(n)$. For each of them, the leader solves a single-level Selection Problem on ${\mathcal{E}_{l}}$ with capacity ${b_{l}}$ in order to determine her solution, and, in addition, solves the adversary’s problem for this fixed leader’s choice. The former can be done in time $O(n)$ and the latter in time at most $R(I)$. The best solution computed in this way clearly gives an optimal leader’s solution. Since the running time $R(I)$ for solving the adversary’s problem cannot be expected to be faster than $O(n)$, which is the time required to solve the follower’s problem for a given scenario, the resulting running time can be written as $O(nR(I))$. ∎
@@
For computing all potential leader’s solutions, the algorithm described in the proof of Theorem 11 requires a running time of $O(n)$ in each of the $O(n)$ iterations, i.e., a total running time of $O(n^{2})$. This part of the algorithm’s total running time can be reduced to $O(n\log n)$ by presorting the leader’s items once in the beginning of the algorithm, as in Algorithm 1; see also Remark 3. However, this does not improve the stated total running time of $O(nR(I))$. Moreover, when applying Theorem 11 for specific types of uncertainty sets below, we do not only use it as a black box and improve on this running time anyway, by doing similar preprocessings that depend on the type of uncertainty set and enable to solve the adversary’s problems faster.
@@ 5.2.1 NP-Hardness
The uncertainty set $\mathcal{U}=\{d_{1},\dots,d_{m+1}\}$ consists of a scenario $d_{i}$ for every edge $e_{i}\in E$ of $G$ and one additional scenario $d_{m+1}$. Instead of defining the scenarios, i.e., follower’s objective functions, $d_{i}$ explicitly, it suffices to give orders of the items that are used in the follower’s greedy algorithm in the respective scenarios. It is always possible to define a function $d_{i}$ for which this order is the unique optimal one for the follower and, in particular, to compute such a function $d_{i}$ in polynomial time; see also Remark 1. Note that the follower never selects more than $b=n+1$ items, so that we can neglect the items after the first $n+1$ ones in the follower’s greedy order. For any $i\in\{1,\dots,m\}$, the scenario $d_{i}$, corresponding to the edge $e_{i}=\{v_{j},v_{k}\}\in E$ with $j<k$, is defined by the order $h_{n+1},h_{n+2},v_{1},\dots,v_{j-1},v_{j+1},\dots,v_{k-1},v_{k+1},\dots,v_{n},h_{n+3}$. The scenario $d_{m+1}$ is defined by the order $h_{n+3},h_{1},\dots,h_{n}$.
@@
Now we turn to leader’s solutions $X\subseteq{\mathcal{E}_{l}}$ that are vertex covers. By definition, for every edge $e_{i}\in E$, at least one of its endpoints is in $X$. Accordingly, for every scenario $d_{i}$, at least one item that is not among the first $n+1$ items of the follower’s greedy order in this scenario is selected by the leader. When the follower now greedily completes the leader’s solution $X$ to $n+1$ items in total, he will never select the item $h_{n+3}$, which is the $(n+1)$-th item in his greedy order. In fact, we get $X\cup Y\subseteq V\cup\{h_{n+1},h_{n+2}\}$. Thus, in each of the scenarios $d_{1},\dots,d_{m}$, the leader’s objective value is $n+1$. In scenario $d_{m+1}$, the follower always selects item $h_{n+3}$ and possibly some items with leader’s cost 00. For the leader’s solution $X$ and the corresponding follower’s solution $Y$, the leader’s objective value is thus $c(X\cup Y)=\lvert X\rvert+n$. This implies that the adversary can be assumed to always select scenario $d_{m+1}$ when the leader has selected a vertex cover $X$, and that, as a leader’s solution, any vertex cover is better than any set that is not a vertex cover. Finally, optimizing the leader’s objective value among the solutions that are vertex covers is equivalent to finding a vertex cover $X\subseteq V$ of minimal size $\lvert X\rvert$. ∎
@@ 5.2.2 Approximation Algorithms
The problem of 2-approximating the Robust Bilevel Selection Problem, with any type of uncertainty set, can be solved in time $O(nR(I))$, where, for an instance $I$ of the Robust Bilevel Selection Problem, $R(I)$ is an upper bound on the running time for solving the corresponding adversary’s problem with any fixed leader’s solution given.
@@ 5.2.3 Exact Algorithms
The Robust Bilevel Selection Problem, with any type of uncertainty set, can be solved in time $O(\min\{2^{n_{l}},{n_{l}^{b}}\}R(I))$, where, for an instance $I$ of the Robust Bilevel Selection Problem, $R(I)$ is an upper bound on the running time for solving the corresponding adversary’s problem with any fixed leader’s solution given.
@@
The problem that has to be solved for a fixed prefix $\overline{Y}\subseteq{\mathcal{E}_{f}}$ is easy to solve: Clearly, the leader must select $b-\lvert\overline{Y}\rvert$ items from ${\mathcal{E}_{l}}\setminus\overline{Y}$. If this set does not contain sufficiently many items, then there is no feasible solution of the desired structure corresponding to the prefix $\overline{Y}$. In addition, the leader may select any number of items from ${\mathcal{E}_{l}}\cap\overline{Y}$, but this has no influence on the overall solution $X\cup Y$ because the follower’s optimal response is always $Y=\overline{Y}\setminus X$. The selection of $b-\lvert\overline{Y}\rvert$ items from ${\mathcal{E}_{l}}\setminus\overline{Y}$ can be made in a greedy way according to the leader’s objective $c$ because there is no difference in the follower’s reaction among different such selections. Therefore, each iteration in the above algorithm boils down to a single-level Selection Problem.

### v3

@@ 1 Introduction
The combination of bilevel optimization and uncertainty is a topic that receives increasing attention currently. In a practical setting in which leader and follower are actually two distinct decision makers, it seems to be very natural to adopt some concept of optimization under uncertainty and assume that, e.g., the leader is uncertain about some parameters of the follower’s problem. A very recent and comprehensive overview of the area of bilevel optimization under uncertainty is given by Beck, Ljubić, and Schmidt in the survey [beckljubicschmidt2023]. It covers both stochastic and robust approaches to model uncertainty, and it classifies various ways of where and how uncertainty can come into play in a bilevel optimization problem. In this article, we focus on uncertainty in the parameters of the follower’s objective function from the leader’s perspective. The situation can be illustrated by considering three decision makers: The leader first chooses a solution, then an adversary (of the leader) chooses an objective function for the follower, and the follower finally computes an optimal solution according to this objective function and depending on the leader’s choice. The leader’s aim is to optimize her own objective value, which depends on the follower’s choice, while the adversary has the opposite objective. Note that defender-attacker-defender games present a similar three-level structure as these robust bilevel optimization problems; see, e.g., [brownetal2006]. However, there is only one objective function in these problems, as the same decision maker acts on the first and on the third level. Moreover, the attack usually does not affect the objective function, but a combinatorial structure concerning the constraints.
@@
A related robust bilevel optimization problem, namely the Robust Bilevel Continuous Knapsack Problem, together with its complexity for different types of uncertainty sets, has been studied in [buchheimhenke2022]. A part of the current work generalizes and extends the results of [buchheimhenke2022], with a stronger emphasis on combinatorial underlying problems. The (Continuous) Bilevel Selection Problem with disjoint item sets is more general than the Bilevel Continuous Knapsack Problem in one aspect, but structurally simpler than it in another; see Section 2.2 for more details. For most of the robust settings we consider, we obtain similar results as in [buchheimhenke2022], but we will also see some differences. In particular, solving the Robust Continuous Bilevel Selection Problem under discrete uncorrelated uncertainty requires some new ideas, which also lead to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem; see Section 6.2.3 and Appendix A. The latter result answers an open question stated in [buchheimhenke2022].
@@
This article is organized as follows. In Section 2, we introduce the Bilevel Selection Problem and discuss several variants of the problem, some of which are distinguished later on. We show how the Bilevel Selection Problem can be solved in polynomial time in Section 2.3. The robust problem version is introduced in Section 3. We then present algorithms for the adversary’s problem, for different types of uncertainty sets, in Section 4, before turning to the robust leader’s problem in Section 5. Here we give polynomial-time algorithms for the simpler case of disjoint item sets in Section 5.1 and show our hardness and algorithmic results for the general case in Section 5.2. Moreover, we give some insights regarding problem variants with continuous variables in Section 6. This includes an example of a bilevel optimization problem whose optimal value may not be attained (Example 25) and some structural properties in Section 6.1 and in the beginning of Section 6.2. We then study the Robust Continuous Bilevel Selection Problem for different types of uncertainty sets. Section 7 concludes. In Appendix A, a new approach developed in Section 6.2 is finally extended to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty.
@@ 2.1 Problem Definition
Note that we do not require the sets ${\mathcal{E}_{l}}$ and ${\mathcal{E}_{f}}$ to be disjoint in general. However, assuming the sets to be disjoint leads to an important special case, actually simplifying the problem in some sense. We denote the cardinalities of the item sets by ${n_{l}}=\lvert{\mathcal{E}_{l}}\rvert$, ${n_{f}}=\lvert{\mathcal{E}_{f}}\rvert$, and $n=\lvert{\mathcal{E}_{l}}\cup{\mathcal{E}_{f}}\rvert$ throughout this article.
@@
In fact, from the leader’s perspective, only this order of the follower’s items is relevant, but not the precise follower’s item costs. Accordingly, when constructing an instance, we will sometimes only define a follower’s greedy order instead of a function $d$; see, e.g., Theorem 15 and Examples 19 and 20. Observe that, given a follower’s greedy order, an appropriate function $d$ (with polynomial-size values) can easily be constructed in polynomial time, e.g., by setting its values to subsequent natural numbers in this order.
@@ 2.2 Problem Variants and Relation to Knapsack Problems
In this section, we introduce and discuss several assumptions and modifications that can be made for the Bilevel Selection Problem as we defined it in Section 2.1. This provides context on related problems in the literature and motivates which problem variants will be distinguished in the rest of this article.
@@ 2.2.1 Item Sets
As mentioned in Section 2.1, we will mainly distinguish the special case of the Bilevel Selection Problem where ${\mathcal{E}_{l}}\cap{\mathcal{E}_{f}}=\emptyset$ from the general one (see Section 5). This special case of the Bilevel Selection Problem can be considered to be a special case of the Bilevel Knapsack Problem that has been studied in [mansietal2012, capraraetal2014]. In fact, the polynomial-time algorithm presented for the Bilevel Selection Problem in Section 2.3, in case of disjoint item sets, can be seen as a special case of the pseudopolynomial-time approaches in [mansietal2012, capraraetal2014]. However, our algorithm also works for the general Bilevel Selection Problem, where the situation is more involved.
Moreover, one could think of a simpler bilevel variant of the Selection Problem in which only the follower selects items in the sense of a single-level Selection Problem, while the leader just decides on the number $b$ of items to select and gets some value (which might be positive or negative) from each selected item. This problem variant can be seen as a special case of the Bilevel Knapsack Problem in which the leader provides the capacity for the follower’s knapsack problem. This problem has been studied in [demperichter2000, capraraetal2014], and the continuous version of it is the underlying Bilevel Continuous Knapsack Problem in [buchheimhenke2022, buchheimhenkeirmai2022]. In fact, this bilevel selection problem is equivalent to a special case of the Bilevel Selection Problem as defined above, with disjoint item sets: Consider ${\mathcal{E}_{l}}$ and ${\mathcal{E}_{f}}$ to be disjoint and the leader’s costs of all leader’s items to be equal to zero. Then, by selecting some subset $X\subseteq{\mathcal{E}_{l}}$ of her items, the leader just determines the capacity $b-\lvert X\rvert$ for the follower and pays her item costs of all items selected by the follower. Therefore, the algorithmic results we obtain in Sections 2.3, 5.1, and 6 can also be applied to this bilevel selection problem, and in fact are similar to the ones in [buchheimhenke2022]. In Appendix A, we extend our new algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncorrelated uncertainty to the Robust Bilevel Continuous Knapsack Problem.
@@ 2.2.3 Maximization
The Selection Problem and also the Bilevel Selection Problem are defined as minimization problems according to Section 2.1, but could analogously be defined as maximization problems, i.e., the decision makers could maximize the total value of their selection of items instead of minimizing the total costs. Then the problems can be more directly seen as special cases of the classical Knapsack Problem and a bilevel variant of it; see also Section 2.2.1.
@@ 5.1 The Special Case of Disjoint Item Sets
Consider the Robust Bilevel Selection Problem with ${\mathcal{E}_{l}}\cap{\mathcal{E}_{f}}=\emptyset$, and with any type of uncertainty set. Suppose that the corresponding adversary’s problem can be solved in time at most $A(I)$, given an instance $I$ of the Robust Bilevel Selection Problem [^2] Note that parts of the structure of the instance $I$ as well as the function $A$ may depend on the type of uncertainty set. together with any leader’s solution. Then the Robust Bilevel Selection Problem, given an instance $I$ with $n$ items, can be solved in time $O(nA(I))$.
@@
The leader enumerates all feasible values of ${b_{l}}$ that determine the capacity used by herself. These are clearly the integer values in the same range as in Algorithm 1, and their number is $O(n)$. For each of them, the leader solves a single-level Selection Problem on ${\mathcal{E}_{l}}$ with capacity ${b_{l}}$ in order to determine her solution, and, in addition, solves the adversary’s problem for this fixed leader’s choice. The former can be done in time $O(n)$ and the latter in time at most $A(I)$. The best solution computed in this way clearly gives an optimal leader’s solution. Since the running time $A(I)$ for solving the adversary’s problem cannot be expected to be faster than $O(n)$, which is the time required to solve the follower’s problem for a given scenario, the resulting running time can be written as $O(nA(I))$. ∎
@@
For computing all potential leader’s solutions, the algorithm described in the proof of Theorem 11 requires a running time of $O(n)$ in each of the $O(n)$ iterations, i.e., a total running time of $O(n^{2})$. This part of the algorithm’s total running time can be reduced to $O(n\log n)$ by presorting the leader’s items once in the beginning of the algorithm, as in Algorithm 1; see also Remark 3. However, this does not improve the stated total running time of $O(nA(I))$. Moreover, when applying Theorem 11 for specific types of uncertainty sets below, we do not only use it as a black box and improve on this running time anyway, by doing similar preprocessings that depend on the type of uncertainty set and enable to solve the adversary’s problems faster.
@@ 5.2.1 NP-Hardness
The uncertainty set $\mathcal{U}=\{d_{1},\dots,d_{m+1}\}$ consists of a scenario $d_{i}$ for every edge $e_{i}\in E$ of $G$ and one additional scenario $d_{m+1}$. Instead of defining the scenarios, i.e., follower’s objective functions, $d_{i}$ explicitly, it suffices to give orders of the items that are used in the follower’s greedy algorithm in the respective scenarios. It is always possible to define a function $d_{i}$ for which this order is the unique optimal one for the follower and, in particular, to compute such a function $d_{i}$ with polynomial-size values in polynomial time; see also Remark 1. Note that the follower never selects more than $b=n+1$ items, so that we can neglect the items after the first $n+1$ ones in the follower’s greedy order. For any $i\in\{1,\dots,m\}$, the scenario $d_{i}$, corresponding to the edge $e_{i}=\{v_{j},v_{k}\}\in E$ with $j<k$, is defined by the order $h_{n+1},h_{n+2},v_{1},\dots,v_{j-1},v_{j+1},\dots,v_{k-1},v_{k+1},\dots,v_{n},h_{n+3}$. The scenario $d_{m+1}$ is defined by the order $h_{n+3},h_{1},\dots,h_{n}$.
@@
Now we turn to leader’s solutions $X\subseteq{\mathcal{E}_{l}}$ that are vertex covers. By definition, for every edge $e_{i}\in E$, at least one of its endpoints is in $X$. Accordingly, for every scenario $d_{i}$, at least one item that is not among the first $n+1$ items of the follower’s greedy order in this scenario is selected by the leader. When the follower now greedily completes the leader’s solution $X$ to $n+1$ items in total, he will never select the item $h_{n+3}$, which is the $(n+1)$-th item in his greedy order. In fact, we get $X\cup Y\subseteq V\cup\{h_{n+1},h_{n+2}\}$. Thus, in each of the scenarios $d_{1},\dots,d_{m}$, the leader’s objective value is $n+1$. In scenario $d_{m+1}$, the follower always selects item $h_{n+3}$ and possibly some items with leader’s cost 00. For the leader’s solution $X$ and the corresponding follower’s solution $Y$, the leader’s objective value is thus $c(X\cup Y)=\lvert X\rvert+n$. This implies that the adversary can be assumed to always select scenario $d_{m+1}$ when the leader has selected a vertex cover $X$, and that, as a leader’s solution, any vertex cover is better than any set that is not a vertex cover. Finally, optimizing the leader’s objective value among the solutions that are vertex covers is equivalent to finding a vertex cover $X\subseteq V$ of minimal size $\lvert X\rvert$.
Note that we have indeed shown strong NP-hardness, since the Vertex Cover Problem does not have any numerical parameters and all values of the constructed functions $c$ and $d_{i}\in\mathcal{U}$ have polynomial size. ∎
@@ 5.2.2 Approximation Algorithms
Consider the Robust Bilevel Selection Problem, with any type of uncertainty set. Suppose that the corresponding adversary’s problem can be solved in time at most $A(I)$, given an instance $I$ of the Robust Bilevel Selection Problem together with any leader’s solution. Then the problem of 2-approximating the Robust Bilevel Selection Problem, given an instance $I$ with $n$ items, can be solved in time $O(nA(I))$.
@@ 5.2.3 Exact Algorithms
Consider the Robust Bilevel Selection Problem, with any type of uncertainty set. Suppose that the corresponding adversary’s problem can be solved in time at most $A(I)$, given an instance $I$ of the Robust Bilevel Selection Problem together with any leader’s solution. Then the Robust Bilevel Selection Problem, given an instance $I$ with ${n_{l}}$ leader’s items and capacity $b$, can be solved in time $O(\min\{2^{n_{l}},{n_{l}^{b}}\}A(I))$.
@@
The subproblem for a fixed prefix $\overline{Y}\subseteq{\mathcal{E}_{f}}$ is easy to solve: Clearly, the leader must select $b-\lvert\overline{Y}\rvert$ items from ${\mathcal{E}_{l}}\setminus\overline{Y}$. If this set does not contain sufficiently many items, then there is no feasible solution of the desired structure corresponding to the prefix $\overline{Y}$. In addition, the leader may select any number of items from ${\mathcal{E}_{l}}\cap\overline{Y}$, but this has no influence on the overall solution $X\cup Y$ because the follower’s optimal response is always $Y=\overline{Y}\setminus X$. The selection of $b-\lvert\overline{Y}\rvert$ items from ${\mathcal{E}_{l}}\setminus\overline{Y}$ can be made in a greedy way according to the leader’s objective $c$ because there is no difference in the follower’s reaction among different such selections. Therefore, each iteration in the above algorithm boils down to a single-level Selection Problem.